NAOJ GW Elog Logbook 3.2

I have some questions.
What is the meaning of the fit in the histograms? Do you have a reason to fit with specific distributions/densities or do you want to find a systematic pattern?

Marc and Yuhang
We have also checked one week data (from 2021/04/25/3pm to 2021/05/01/3pm)
We made two plots, the first plot includes all the data in this week. But the second plot excluded the last part of the data. This is because an earthquake might happen around the end of the one week data. This makes we see a peak in the first plot.
To compare different mirrors angular drift, we calculated RMS value of data and summarized them as the attached table. [unit: urad]
|
PR pit |
PR yaw |
BS pit |
BS yaw |
INPUT pit |
INPUT yaw |
END pit |
END yaw |
angular drift RMS (including earthquake) |
59 |
4.5 |
14.35 |
8.85 |
31.24 |
15.81 |
9.81 |
2.86 |
angular drift RMS earthquake excluded |
62.46 |
3.98 |
10.39 |
9.07 |
29.42 |
15.49 |
9.08 |
2.66 |
From this table, we see PR pitch moved the most. But INPUT pitch also moved a lot. However, END pitch didn't move as large as PR and INPUT pitch. Since PR, INPUT and END have the same configuration, their difference in pitch drift indicates that this is not a design problem.
We know that temperature change influences mirrors drift. But PR, BS, and INPUT are all in the TAMA central room, they should have the same temperature change. But they didn't show the same drift. Form the airconditioner location, PR chamber is almost facing an air conditioner, which may cause temperature related problem. To confirm this is not the fault of air conditioner, we plan to switch off it for one or two days during weekends to check.

According to the OPLEV signal calibration value in elog1874, I plotted the four suspended mirrors drift during a time scale of one day(from 2021/05/12 1pm to 2021/05/13 1pm JST).
In figure 1, all four suspended mirrors are compared. It is very clear that PR pitch has a far larger angular drift compared with other mirrors.
The PR mirror changed by 100urad pk-pk during one day. Considering the distance from PR to END mirrors, the 1e-4*300 = 3e-2 m = 3cm.
Other mirrors other DOF changed by 10urad pk-pk during one day, which is ten times smaller than PR pitch.
Considering this measurement result, I think we should make beam pointing loop act on PR mirror.
Marc and Yuhang
We have also checked one week data (from 2021/04/25/3pm to 2021/05/01/3pm)
We made two plots, the first plot includes all the data in this week. But the second plot excluded the last part of the data. This is because an earthquake might happen around the end of the one week data. This makes we see a peak in the first plot.
To compare different mirrors angular drift, we calculated RMS value of data and summarized them as the attached table. [unit: urad]
|
PR pit |
PR yaw |
BS pit |
BS yaw |
INPUT pit |
INPUT yaw |
END pit |
END yaw |
angular drift RMS (including earthquake) |
59 |
4.5 |
14.35 |
8.85 |
31.24 |
15.81 |
9.81 |
2.86 |
angular drift RMS earthquake excluded |
62.46 |
3.98 |
10.39 |
9.07 |
29.42 |
15.49 |
9.08 |
2.66 |
From this table, we see PR pitch moved the most. But INPUT pitch also moved a lot. However, END pitch didn't move as large as PR and INPUT pitch. Since PR, INPUT and END have the same configuration, their difference in pitch drift indicates that this is not a design problem.
We know that temperature change influences mirrors drift. But PR, BS, and INPUT are all in the TAMA central room, they should have the same temperature change. But they didn't show the same drift. Form the airconditioner location, PR chamber is almost facing an air conditioner, which may cause temperature related problem. To confirm this is not the fault of air conditioner, we plan to switch off it for one or two days during weekends to check.

Abe, Marc
Following the 3 maps measurements we performed again the bulk calibration where this time we moved the imagining unit by 0.32 mm in order to compensate the thickness difference between surface and bulk reference samples. We got the following result (also see last figure of this entry) :
AC_bulkref = 0.0731;
DC_bulkref = 4.164;
ACDC = 0.01755;
P_in = 29.5e-3;
P_t = 13.3e-3;
T_bulkref = P_t/P_in;
abs_bulkref = 1.04;
R_bulk = AC_bulkref/(DC_bulkref*sqrt(T_bulkref)*P_in*abs_bulkref) = 0.852 cm/W
Using this new calibration factor and using :
P_t = 6.25 W;
P_in = 7.322 W;
the absorption of this sample seems to be around 60 ppm/ cm for all 3 maps (see attached figures)
Today we will double check the bulk calibration as the change was quite larger than expected.

Abe, Marc
We modified the analysis to better estimate the mean and standard-deviation of absorption measurements.
The corrected results are attached to this entry.
Today I will remove the first contact that we applied on this sample and cross-checked if it affected the absorption measurement.

[Aritomi, Yuhang]
First we measured nonlinear gain again. We measured BAB transmission from OPO. When we used 40mW green, we decreased CC1 gain from 2 to 1.
green power (mW) | BAB maximum (V) | OPO temperature (kOhm) | p pol PLL (MHz) |
0 | 0.0552 | 7.194 | 305 |
40.9 | 1.02 | 7.194 | 195 |
The measured nonlinear gain is 18.5 with 40.9 mW green while the theoretical value should be g = 1/(1-sqrt(40.9/80))^2 = 12.3. We don't know why they are different.
We found an oscillation in FC lock with green injection of 27mW and FC gain of 1.3. The green injection power was larger so we decreased FC gain from 1.3 to 1.
We measured CCFC with 40.9 mW pump green and 50% pick off. Fig 1 shows the CCFC error signal with some demodulation phase and Fig 2 shows CCFC calibration signal when CCSB are off resonance and CC1 is scanned. The amplitude of CCFC calibration signal is 524mVpp which is 10 times larger than before.
We will try with 20% pick off and thorlabs BSS11 seems good for BS (20% reflection for p pol).

The measurement performed during the golden week were using too low power (I_laser =3A ie P_laser~3.5W) that made it hard to really acquire signals.
The measurements have been restarted with higher power (I_laser=6A ie P_laser~8W?).
We'll have to check carefully the incident and transmitted power after the measurements.
Abe, Marc
Following the 3 maps measurements we performed again the bulk calibration where this time we moved the imagining unit by 0.32 mm in order to compensate the thickness difference between surface and bulk reference samples. We got the following result (also see last figure of this entry) :
AC_bulkref = 0.0731;
DC_bulkref = 4.164;
ACDC = 0.01755;
P_in = 29.5e-3;
P_t = 13.3e-3;
T_bulkref = P_t/P_in;
abs_bulkref = 1.04;
R_bulk = AC_bulkref/(DC_bulkref*sqrt(T_bulkref)*P_in*abs_bulkref) = 0.852 cm/W
Using this new calibration factor and using :
P_t = 6.25 W;
P_in = 7.322 W;
the absorption of this sample seems to be around 60 ppm/ cm for all 3 maps (see attached figures)
Today we will double check the bulk calibration as the change was quite larger than expected.
It seems that the bulk calibration was overestimated. This is especially apparent when computing its transmission that was 45% instead of the expected 55%.
I performed again the bulk calibration and got :
AC_bulkref = 0.062;
DC_bulkref = 4.14;
P_in = 26.4e-3;
P_t = 13.1e-3;
T_bulkref = P_t/P_in
abs_bulkref = 1.04;
R_bulk = AC_bulkref/(DC_bulkref*sqrt(T_bulkref)*P_in*abs_bulkref) = 0.7743 W/cm
I used this new calibration to compute again the absorption map of the sample (see the 3 attached figures.
In the figure, the absorption is extracted from a fit using 2 normal distributions.
Here I also add the overall mean and standard deviation of each map (ie without any fittting) :
XY | YZ | XZ | |
mean [ppm] | 70 | 50 | 48 |
std [ppm] | 8 | 35 | 37 |

Today I putted first contact on one side of the SHINKOSHA evaluation plate.
We'll remove and clean the other side tomorrow.

Thank you.
I have updated the absorption value using the bulk calibration measured just after removing this sample (see elog 2480) and the corrected absorption is 302 +/- 108 ppm.

Yuhang and Michael
We finished setting up the modulators.
Set up and aligned FI: 5.15 mW transmission, 5.54 mW incident (92.96% tranmission vs Thorlabs 92% specification)
Set up HWP and PBS to check vertical polarisation required for EOM: 5.17 mW transmit, 0.03 mW reflect, 5.35 mW incident (PBS), 5.97 mW incident (HWP)
Set up EOM: 5.63 mW transmission, 6.04 incident.
At this point the power meter was seen to be not particularly reliable for alignment, since it seems to register too much power at the edge of the detector compared to the center. The EOM was centered by using the IR detector card and seeing that the beam passes through the center. Then, we added AOM.
Afterwards, we measured the beam profile coming out of the modulators. In 2421, the beam profile was measured coming from the lens that was previously fixed on the bench, as indicated in that entry. From that measurement I used a beam waist of 150 µm with position z = 0 defined by the position of the last preset lens. However, that beam waist was an average of the horizontal and vertical beam waist measurements, which were quite different. But perhaps that could have also been due to the alignment of the beam on the beam profiler, which in 2425 was seen to be off centered in the measurements close to z = 0. Using Jammt with Thorlabs lens database, an f = 100 mm lens placed at z = 0.125 m gives a waist of 218 µm at z = 0.272 (figure 1). However, as mentioned in 2482, the lenses shown were slightly adjusted in order to properly level the beam at 76 mm height. So the prediction may not be completely accurate anymore.
The measured beam profile is shown in figure 2. The predicted beam seems to be a bit too far forward and too large.
Readjusting the mode matching telescope with a waist of 193 µm at 13.3 cm with respect to the f = 100 mm lens gives the result shown in figure 3. The spacing between the two modematching lenses (f = 40 mm and f = 75 mm), as well as the space from the f = 75 mm to the OPO waist should still be sufficient.

Yuhang and Michael
We just did some adjustments to the beam using the lenses shown, to make sure the beam is travelling at about 76mm height along the screw hole line. It rises by about 1mm over ~50cm of travel. The mounts have also been adjusted to match the beam height.
This work will be on hold while the KAGRA ETMY is being cleaned.

Nice results!
It seems that Shinkosha still has its main feature: the prominent spider-web absorption.
But the absorption is really high this time.

Abe, Aritomi, Marc
Yesterday we removed the shinkosha evaluation plate and its holder from the translation stage.
We installed back the bulk reference sample and got R_bulk ~0.773 W/cm with
AC_bulkref = 0.0555 V;
DC_bulkref = 4.103 V;
P_in = 24.2e-3 W;
P_t = 11.7e-3 W;
abs_bulkref = 1.04;
Then, we brought the KASI sample inside the clean room and cleaned it with alcool while checking with the green light.
With this sample, the IU is now at z=43.1mm and the mirror centers are :
X_center = 327.5 mm
Y_center = 122.95 mm
Z_center = 62.5 mm (we had to increase the laser power to more than 1 W (~3W) to clearly see the phase change at the sample surfaces)
We added new X limit (lower vertical) on Zaber and started a circular map at Z_center with 2 cm radius.
Our initial guess of the absorption is abs~78 ppm/cm given by
AC = 2.3e-4 V;
DC = 4.242 V;
P_in = 3.24 W;
P_t = 2.77 W;
T_sample = P_t/P_in;
R_bulk = 0.7730 W/cm;
mat_correction=3.34;
abs = AC/(DC*P_in*sqrt(T_sample)*R_bulk)*1e6*mat_correction

Yuhang and Michael
We did some more inspection of the OPO replacement setup at ATC. Perhaps we will need some more low f lenses.
I recalculated the error signal and mode matching shown in 2469.
We can obtain a much larger error signal in reflection when the meniscus is used as the input mirror. Even with <10 mW incident power it should not be a problem. Here I use 4 mW incidence, 40 MHz modulation and 0.3 modulation depth to achieve ~160 mVpk error signal in the linear range. Figure 1 shows the error signals and power outside of the cavity. Figure 2 shows the scan of demodulation phase versus reflection and transmission photodetector signals. Figure 3 shows the circulating power of 160 mW, for 4 mW input.
I also show a beam profile with a more complete mode matching telescope into the OPO. I did a mode matching telescope calculation using JamMT and the database of Thorlabs lenses. We wish to obtain a beam waist of 20.66 µm within the OPO cavity using two lenses. Judging by the OPO setup in TAMA, it would be good to leave approximately 15cm OPO to lens and lens to lens.
In 2469 I used an f = 100mm lens placed at 175mm from the last preset optic on the ATC bench. The beam waist of 150µm is located within 1cm of the last preset optic on the ATC bench This creates another 150µm beam waist 175mm from the lens. The rationale for doing this was:
i) to have a couple of steering mirrors before this lens, so that the beam would be level at 76mm height when going through the FI/EOM/AOM. However, Yuhang suggested to simply move the 100mm lens to achieve beam levelling.
ii) to reduce the size of the beam waist, and make the beam less than 1/5th of the EOM aperture size for a reasonable distance.
This time I used an f = 100mm lens placed 125mm from the last preset optic (as per photograph and measurement in 2439). The mode matching telescope, calculated by JamMT, is shown in figure 4. We should have enough space for the modulators between the f = 100mm and f = 40mm (90mm FI, HWP, 56mm EOM, 22mm AOM). Steering mirrors can fit in the space between the f = 40mm and f = 75mm, and then there should be a beam splitter between the f = 75mm and the OPO cavity.

Marc, Matteo, Yuhang
We used the green light to check possible dust on the mirror and took a picture after already using the air dust once.
The mirror is really dirty...
We then used again the air dust and could remove most of the large dust particles but not the smallest one (not visible on picture so no 'after picture').
We can still see the ethanol/acetone trace at the mirror edge so the mirror got dirty either during shipping or during the few weeks it stayed inside the clean room without cover.
We will clean it with first contact after the golden week.

Before the measurement we also moved the IU to z=68.6mm.
I'll add to the wiki the IU position for surface and bulk reference as well as TAMA size, 1cm shinkosha and KAGRA samples.

Abe, Marc, Matteo
Following the measurement of the absorption map at z=51.4 mm, we performed 2 others map in the XZ and YZ planes.
They correspond respectively to figure 1 and 2.
An important point to notice is that there seems to be a point absorber on the surface that distorts the absorption scale (up to 4000 ).
We decided to turn off the ir laser just in case.
Anyway looking at the histogram, it seems that the nominal absorption is around 400 ppm/cm (more precise analysis to come).
In figure 3 you can see a combination of the 3 maps where the colorbar scale has been limited to the maximum of the circular map (ie 1500 ppm/cm).
Marc, Matteo, Yuhang
We used the green light to check possible dust on the mirror and took a picture after already using the air dust once.
The mirror is really dirty...
We then used again the air dust and could remove most of the large dust particles but not the smallest one (not visible on picture so no 'after picture').
We can still see the ethanol/acetone trace at the mirror edge so the mirror got dirty either during shipping or during the few weeks it stayed inside the clean room without cover.
We will clean it with first contact after the golden week.
Nice results!
It seems that Shinkosha still has its main feature: the prominent spider-web absorption.
But the absorption is really high this time.
Thank you.
I have updated the absorption value using the bulk calibration measured just after removing this sample (see elog 2480) and the corrected absorption is 302 +/- 108 ppm.
Abe, Marc
We modified the analysis to better estimate the mean and standard-deviation of absorption measurements.
The corrected results are attached to this entry.
Today I will remove the first contact that we applied on this sample and cross-checked if it affected the absorption measurement.
I have some questions.
What is the meaning of the fit in the histograms? Do you have a reason to fit with specific distributions/densities or do you want to find a systematic pattern?
The absorption distribution is fitted with 2 normal distributions.
I thought it could be useful for the case of XZ and YZ maps (where there are measurement points outside the sample) because it allows to remove the effects of absorption outside the sample and point defects/dust on the surface.
But I agree that it might not be the most suitable distribution, especially for the shinkosha samples...

The absorption map and absorption distribution of the 1 cm thick shinkosha sample is attached to this entry.
It has a mean absorption of 355 +/- 127 ppm/cm

Abe, Marc, Matteo
This morning, we installed the 1cm thick SHINKOSHA sample.
We recentered the beam dump on the laser unit and covered back this unit.
Then, we increased the power to ~300 mW to check the sample center and inclinaison.
We checked its X,Y and Z centers :
X center = 398.195 mm
Y center = 121.5 mm
Z center = 51.4 mm
Then, we checked the z center at the left, right, bottom and top position of the sample with extremal positions separated by 13 cm.
We found a 0.8 mm vertical tilt of the mirror and a 0.4 mm horizontal tilt.
Finally, we increased the laser power to 3.46W and started the measurement of absorption map of this sample.
We also did a preliminary evaluation of the absorption as
AC = 3.6e-3 V;
DC = 4.71 V;
P_in = 3.46 W;
P_t = 2.94 W;
T_sample = P_t/P_in;
R_bulk =0.7414 W/cm;
mat_correction=3.34;
abs = AC/(DC*P_in*sqrt(T_sample)*R_bulk)*1e6*mat_correction ~1080 ppm/cm
Before the measurement we also moved the IU to z=68.6mm.
I'll add to the wiki the IU position for surface and bulk reference as well as TAMA size, 1cm shinkosha and KAGRA samples.

Yuhang and Michael
We took some equipment from TAMA - camera + TV setup for when we will assemble the OPO cavity there, as well as some mounts to keep the beam height correct (should be 76mm).
The Faraday Isolator was placed and characterised. There was a scratch on the polariser window which reduced transmission somewhat, so the fine tuning of the translation stage on the FI mount was adjusted to avoid this.
The power of the Lightwave laser placed there does not seem very stable and drifts from minute to minute.
If the Faraday isolator is not well aligned along the beam axis, the beam becomes astigmatic.
Using the Thorlabs power meter, we measure the following:
- Initial transmission adjustment: 4.79 mW transmit, 5.03 mW incident
- Blocking adjustment: 0.25 µW transmit, 5.185 mW incident
- Placed back in transmitting mode: 4.991 mW transmit, 5.343 mW incident
This gives us 93.4% power transmission and 43.2 dB isolation (Thorlabs specifies 92% transmission and 35 dB isolation).